The number of tickets sold each day for an upcoming performance od Handel's Messiahis given by N(x)=-0.6x^2+18x+14, where x is the number of days since the concert first was announced.When will daily tickets peak and how many will be sold that day?
Pls help...ASAP.
thanks

Answer by stanbon(75887)

The number of tickets sold each day for an upcoming performance od Handel's Messiahis given by N(x)=-0.6x^2+18x+14, where x is the number of days since the concert first was announced. When will daily tickets peak and how many will be sold that day?
------
Maximum occurs when x = -b/2a = -18/(2*-0.6) = 15 (the day it peaks)
---
N(15) = -0.6(15)^2+18(15)+14 is the number sold.
N(15) = 149
=========================
cheers,
Stan H.